OpenStudy (anonymous):

@Xmoses1 dude plz help me or anyone

OpenStudy (anonymous):

Which equation represents the direct variation shown in the graph? A graph shows the points 4 comma ten, 6 comma fifteen, and 8 comma twenty with a line crossing all of them and going through the point zero comma zero y equals one point five times x y equals 2 times x y equals two point five times x y equals 5 times x

OpenStudy (xmoses1):

Sorry idk this very well :/

OpenStudy (anonymous):

help me ive been on this forever

OpenStudy (anonymous):

@emmigrace222

OpenStudy (math&ing001):

When you look at all those values you notice that each time 2.05 is added. Ex: 6.15=4.10+2.05 and 8.20=6.15+2.05 What do you think the equation should be?

OpenStudy (anonymous):

idfk

OpenStudy (emmigrace222):

ok im going to read this and see what i can doo

OpenStudy (anonymous):

kkkkkkkkkkkkkkkkk

OpenStudy (emmigrace222):

shoot i have no idea try google\bing

OpenStudy (anonymous):

kkkkk

OpenStudy (anonymous):

di u get it yet